2.1. The following is a recipe for making 18 scones: 1 cup white sugar, cup butter, 2 teaspoons vanilla essence, 1 cups flour, 2 eggs, 1 teaspoons baking powder, cup of milk. On your birthday you decide to use this recipe to make scones for the staff at your school. How would you adjust the recipe so that you can make 60 scones?                                      


1
Expert's answer
2022-04-06T10:23:49-0400

"Every\,\,ingredient\,\,multiplied\,\,by\,\,\frac{60}{18}=\frac{10}{3}\\Thus\,\,we\,\,have\,\,\\\frac{10}{3}\,\,cup\,\,white\,\,sugar\\\frac{10}{3}\,\,cup\,\,butter\\\frac{20}{3}\,\,teaspoons\,\,vanilla\\\frac{10}{3}\,\,cups\,\,floor\\\frac{20}{3}\,\,eggs\\\frac{10}{3}\,\,teaspoons\,\,baking\,\,powder\\\frac{10}{3}\,\,cup\,\,of\,\,milk"